# Non-Stationary Covariance Functions for Spatial Data on Linear Networks

### v1

@@ Abstract
We introduce a novel class of non-stationary covariance functions for random fields on linear networks that allows both the variance and the correlation range of the random field to vary spatially. The proposed covariance functions are useful to model random fields with a spatial dependence that is locally isotropic with respect to the resistance metric, a distance that reflects the topology of the network. We assess the statistical and computational performance of a weighted local likelihood estimator for the proposed models using synthetic data generated on the street network of the University of Chicago neighborhood.
keywords:
Generalized Stieltjes functions , Laplacian matrix , Log Gaussian Cox process , Positive definite functions , Gaussian random field , Resistance metric , Weighted local likelihood
[…]
organization=Department of Statistics, Pontificia Universidad Católica de Chile,addressline=Avenida Vicuña Mackenna 4860, city=Macul, postcode=7820436, state=Santiago, country=Chile
@@ 1 Introduction
Data indexed by locations on a network arise naturally in a variety of applications, including processes on road, river, and electrical networks. Recent developments in the theory and methods of spatial statistics on linear networks can be found in Ver Hoef and Peterson (2010), Baddeley et al. (2017), Anderes et al. (2020), Cronie et al. (2020), Moradi and Mateu (2020), Baddeley et al. (2021), Porcu et al. (2023), Bolin et al. (2024), Møller and Rasmussen (2024), Moradi and Sharifi (2024), González-Pérez et al. (2025) and Alegría et al. (2026).
@@
A second-order Gaussian random field (GRF) $\{X(s):s\in\mathcal{L}\}$ is completely characterized by its mean function, which we assume to be zero for simplicity, and by its covariance function $K(s,t)={\rm cov}[X(s),X(t)]$, $s,t\in\mathcal{L}$. The construction of valid and versatile covariance functions is challenging because of the requirement that the covariance matrix of $(X(s_{1}),\ldots,X(s_{N}))^{\top}$, given by
@@
should be positive semidefinite, for all $N\in\mathbb{N}$ and for all collection of sites $s_{1},\dots,s_{N}\in\mathcal{L}$. In other words, $K$ must be a positive semidefinite function. The spectral theory of random fields on Euclidean spaces is useful to check for positive definiteness of a candidate function (see, e.g., Yaglom, 1987). In the network setting, however, analogous characterizations are unavailable.
@@
with $\text{d}(\cdot,\cdot)$ standing for a distance metric on $\mathcal{L}$. In practice, treating network coordinates as ordinary Euclidean coordinates can distort the proximity between locations and produce unrealistic results. To address this, Anderes et al. (2020) established positive definiteness for certain classes of isotropic covariance functions under the resistance metric (an extension of the resistance metric from electrical network theory) and showed that, under the geodesic (shortest-path) metric, these classes are valid only under restrictive assumptions on the network structure. Both the resistance and geodesic metrics, unlike the Euclidean metric, take the network topology into account.
@@
Although isotropy is a convenient assumption, it can be overly restrictive. In geostatistical applications, there is increasing demand for more flexible models, particularly when analyzing data over large domains where the properties of the process may vary across space. See Bolin et al. (2025) for a recent development of non-stationary random fields on metric graphs formulated as solutions to stochastic partial differential equations, and an application to traffic speed data. In the context of log Gaussian Cox processes (LGCPs) (Cox, 1955), where the intensity of point occurrences is governed by a log Gaussian random field, isotropy of the underlying GRF prevents the number of points and the spatial extent of point clusters from varying across the domain. This limitation has been studied for point processes in Euclidean spaces by Dvořák et al. (2019). Similar issues arise on linear networks, as discussed by Baddeley et al. (2017), encouraging the development of more flexible models.
@@
Motivated by this, we propose a general class of non-stationary covariance functions on linear networks, that is, covariance functions not solely determined by distance. The proposed covariance models allow both the variance and the correlation range of the random field to vary across the network, providing a network analogue of models previously proposed in Euclidean spaces (Paciorek, 2003; Paciorek and Schervish, 2006; Anderes and Stein, 2011; Kleiber and Nychka, 2012; Mateu et al., 2013; Liang and Marcotte, 2016). Our framework produces covariance functions that are locally isotropic with respect to the resistance metric, ensuring interpretability and alignment with the topology of the network.
@@
We exploit the connection between the proposed covariance functions and the class of generalized Stieltjes functions (see Koumandos and Pedersen, 2019; Menegatto, 2020 for a detailed treatment) to provide examples of non-stationary covariance functions in closed form.
@@
To address statistical inference, we revisit the weighted local likelihood approach, originally introduced by Anderes and Stein (2011) for Euclidean spaces, and adapt it to the network setting. We then conduct a simulation study using synthetic data generated on the street network of the University of Chicago neighborhood, in which we estimate the local correlation range, examine the behavior of the method under different weighting schemes, and evaluate its computational performance.
@@
The paper is organized as follows. Section 2 reviews the definition and properties of the resistance metric for linear networks. Section 3 presents the main theorem, examples, and illustrations. Section 4 describes the weighted local likelihood method and reports the results of the simulation study. Section 5 concludes the paper with a discussion and directions for future research.
@@ 2 Resistance metric
This section reviews the definition and properties of the resistance metric introduced in Anderes et al. (2020), as this metric is essential for building the non-stationary covariance functions proposed in this work.
We begin by representing the linear network as a graph $(\mathcal{V},\mathcal{E})$, where the vertex set $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ corresponds to the endpoints of $\mathcal{L}$, and the edge set $\mathcal{E}$ corresponds to the linear segments $\ell_{1},\dots,\ell_{m}$. The length of $\ell_{i}$ is denoted by $|\ell_{i}|$. We write $v_{i}\sim v_{j}$ if the distinct vertices $v_{i}$ and $v_{j}$ are connected by an edge. For any pair of vertices $v_{i},v_{j}\in\mathcal{V}$, the conductance function is defined as
@@
where $v^{*}\in\mathcal{V}$ is an arbitrary reference vertex and $c(v_{i})=\sum_{v\sim v_{i}}\text{con}(v_{i},v)$. The resulting matrix can be viewed as a modification of the graph Laplacian, where the inclusion of an additional unit term on the diagonal guarantees strict positive definiteness. This construction is invariant to the choice of $v^{*}$ (Anderes et al., 2020).
@@
    where $B_{1},\dots,B_{m}$ are independent standard Brownian bridges, each defined on an interval of length $|\ell_{i}|$, starting and ending at zero, and independent of $A_{0}$. Since each $B_{i}$ is defined on an interval, writing $B_{i}(s)$ implicitly identifies points $s$ along $\ell_{i}$ with points in this interval via arc length parameterization (see Møller and Rasmussen, 2024 for further details).
@@
The resistance metric on $\mathcal{L}$ is defined as the variogram of the auxiliary GRF in Equation (1):
$\text{d}_{R}(s,t)={\rm var}[A(s)-A(t)],\qquad s,t\in\mathcal{L}.$ (2)
@@
This metric extends the classical resistance metric from electrical network theory (Klein and Randić, 1993), originally defined on the vertex set, to the continuum of points along the edges of the graph. When evaluated on any additional points along the edges, $\text{d}_{R}$ coincides with the metric obtained from the corresponding discrete electrical network. Interestingly, the resistance and geodesic distances coincide when $\mathcal{L}$ is a tree, that is, a connected graph with no cycles. Møller and Rasmussen (2024) provide more fully developed algebraic expressions for the resistance metric, which facilitate its implementation.
@@
We turn to a description of the range of isotropic covariance functions attainable using the resistance metric. A function $\varphi:[0,\infty)\to\mathbb{R}$ is said to be completely monotone on $0,\infty)$ if it is continuous on $0,\infty)$, infinitely differentiable on $(0,\infty)$, and satisfies $(-1)^{j}\varphi^{(j)}(z)\geq 0$, $z>0$, for every integer $j\geq 0$, where $\varphi^{(j)}$ denotes the $j$th derivative of $\varphi$ and $\varphi^{(0)}=\varphi$. Anderes et al. (2020) showed that for any completely monotone function $\varphi$, the function $K(s,t)=\varphi\big(\text{d}_{R}(s,t)\big)$ is (strictly) positive definite on $\mathcal{L}\times\mathcal{L}$. As a consequence, many popular parametric classes of covariance functions developed for Euclidean spaces are also valid with respect to the resistance metric for general linear networks (see Anderes et al., 2020, Table 1).
@@
In contrast, the geodesic distance provides a more restrictive alternative. For example, if $\mathcal{L}$ contains three distinct paths connecting two points $s,t\in\mathcal{L}$, the classical exponential covariance function, defined with respect to the geodesic metric, is no longer positive definite on this network (Anderes et al., 2020).
@@
In the context of point processes on linear networks, Møller and Rasmussen (2024) emphasize the advantages of the resistance metric over the geodesic metric in terms of modeling flexibility. In particular, the use of $\text{d}_{R}$ allows one to model stronger clustering effects in Cox processes and higher degree of repulsiveness in determinantal point processes.
@@ 3.1 Main result
Let $a:\mathcal{L}\rightarrow(0,\infty)$ and $b:\mathcal{L}\rightarrow(0,\infty)$, and define $\alpha(s,t)=[a(s)+a(t)]/2$. Let $\text{d}_{R}(s,t)$ be the resistance metric between $s$ and $t$, and let $F$ be a positive finite measure on $(0,\infty)$. Then, the function
$K(s,t)=\displaystyle\int_{0}^{\infty}\frac{\sqrt{b(s)b(t)}}{\sqrt{\alpha(s,t)+\text{d}_{R}(s,t)w}}\text{d}F(w),\qquad s,t\in\mathcal{L},$ (3)
@@
Our proof of Theorem 3.1, presented below, adapts ideas introduced in Paciorek (2003) for establishing positive definiteness of non-stationary covariance functions in Euclidean spaces. To the best of our knowledge, the present work is the first to extend this approach to linear networks. The fact that the resistance metric is defined as the variogram of a specific auxiliary GRF gives the proof a distinctive character.
Proof.
A direct calculation shows that, for all $w>0$,
$\frac{1}{\sqrt{\alpha(s,t)+\text{d}_{R}(s,t)w}}$ $=$ $\frac{1}{\sqrt{\alpha(s,t)}}\int_{\mathbb{R}}\exp\left\{-\frac{w\text{d}_{R}(s,t)y^{2}}{2\alpha(s,t)}\right\}\frac{1}{\sqrt{2\pi}}\exp\left\{-\frac{y^{2}}{2}\right\}\text{d}y,$ (4)
where the expectation in (4) is taken with respect to $Y\sim\mathcal{N}(0,1)$. Consequently, $K$ in Equation (3) can be written as
$K(s,t)=\frac{\sqrt{b(s)b(t)}}{\sqrt{\alpha(s,t)}}\int_{0}^{\infty}{\rm E}\left[\exp\left\{-\frac{w\text{d}_{R}(s,t)Y^{2}}{2\alpha(s,t)}\right\}\right]\text{d}F(w),\qquad s,t\in\mathcal{L}.$ (5)
Let $\phi(\cdot;\mu,\sigma^{2})$ denote the probability density function of a normal distribution with mean $\mu$ and variance $\sigma^{2}$. Using a convolution argument, we obtain
$\int_{\mathbb{R}}\phi\left(\xi;\frac{A(s)}{\sqrt{2}},\frac{a(s)}{4w}\right)\phi\left(\xi;\frac{A(t)}{\sqrt{2}},\frac{a(t)}{4w}\right)\text{d}\xi=\frac{\sqrt{w}}{\sqrt{\pi\alpha(s,t)}}\exp\left\{-\frac{w\big[A(s)-A(t)\big]^{2}}{2\alpha(s,t)}\right\},$ (6)
for every $w>0$, with this expression conditioned on the values, $A(s)$ and $A(t)$, of the auxiliary GRF defined in Equation (1).
Notice that $A(s)-A(t)\sim\mathcal{N}(0,\mathrm{d}_{R}(s,t))$ and, consequently, $\big[A(s)-A(t)\big]/\sqrt{\mathrm{d}_{R}(s,t)}$ is equal in distribution to $Y\sim\mathcal{N}(0,1)$. Taking the expectation in (6) with respect to $(A(s),A(t))^{\top}$, we obtain
${\rm E}\left[\int_{\mathbb{R}}\phi\left(\xi;\frac{A(s)}{\sqrt{2}},\frac{a(s)}{4w}\right)\phi\left(\xi;\frac{A(t)}{\sqrt{2}},\frac{a(t)}{4w}\right)\text{d}\xi\right]=\frac{\sqrt{w}}{\sqrt{\pi\alpha(s,t)}}\,{\rm E}\left[\exp\left\{-\frac{w\text{d}_{R}(s,t)Y^{2}}{2\alpha(s,t)}\right\}\right],$ (7)
where the right-hand side of Equation (7) simplifies to an expectation with respect to $Y\sim\mathcal{N}(0,1)$. Therefore, using Equation (7) and Fubini’s theorem, the function $K$ in Equation (5) can be written as
$K(s,t)=\sqrt{\pi}\,{\rm E}\left[\int_{0}^{\infty}\int_{\mathbb{R}}\sqrt{b(s)b(t)}\,\phi\left(\xi;\frac{A(s)}{\sqrt{2}},\frac{a(s)}{4w}\right)\phi\left(\xi;\frac{A(t)}{\sqrt{2}},\frac{a(t)}{4w}\right)\text{d}\xi\,w^{-1/2}\,\text{d}F(w)\right].$
$\sum_{i=1}^{N}\sum_{j=1}^{N}r_{i}r_{j}K(s_{i},s_{j})= \sum_{i=1}^{N}\sum_{j=1}^{N}r_{i}r_{j}\sqrt{\pi}\,{\rm E}\bigg[\int_{0}^{\infty}\int_{\mathbb{R}}\sqrt{b(s_{i})b(s_{j})}\,\phi\bigg(\xi;\frac{A(s_{i})}{\sqrt{2}},\frac{a(s_{i})}{4w}\bigg)\\ \phi\bigg(\xi;\frac{A(s_{j})}{\sqrt{2}},\frac{a(s_{j})}{4w}\bigg)\text{d}\xi\,w^{-1/2}\,\text{d}F(w)\bigg],$
that is,
$\sum_{i=1}^{N}\sum_{j=1}^{N}r_{i}r_{j}K(s_{i},s_{j})=\sqrt{\pi}\,{\rm E}\left[\int_{0}^{\infty}\int_{\mathbb{R}}\left\{\sum_{i=1}^{N}r_{i}\sqrt{b(s_{i})}\,\phi\left(\xi;\frac{A(s_{i})}{\sqrt{2}},\frac{a(s_{i})}{4w}\right)\right\}^{2}\text{d}\xi\,w^{-1/2}\,\text{d}F(w)\right]\geq 0.$
Therefore, $K$ is positive semidefinite.
@@
In this framework, the function $b$ provides a way to modulate the locally varying variance of the field, while the function $a$ regulates how quickly the covariance decreases with distance. In particular, $a$ can be adjusted to influence the locally adaptive correlation range (the distance beyond which correlations fall below $0.05$).

### v2

@@ Abstract
We introduce a novel class of non-stationary covariance functions for random fields on linear networks that allows both the variance and the correlation range of the random field to vary spatially. The proposed covariance functions are useful to model random fields with a spatial dependence that is locally isotropic with respect to the resistance metric, a distance that reflects the topology of the network. The framework admits explicit stochastic representations of the associated random fields and can be naturally extended to matrix-valued covariance functions for vector-valued random fields. We assess the statistical and computational performance of a weighted local likelihood estimator for the proposed models using synthetic data generated on the street network of the University of Chicago neighborhood.
Keywords: Generalized Stieltjes functions; Laplacian matrix; Log Gaussian Cox process; Positive definite functions; Gaussian random field; Resistance metric; Vector-valued random field; Weighted local likelihood
Departamento de Estadística, Pontificia Universidad Católica de Chile,
Santiago, Chile
February 20, 2026
@@ 1 Introduction
Data indexed by locations on a network arise naturally in a variety of applications, such as processes on road, river, and electrical networks. Recent developments in the theory and methodology of spatial statistics on linear networks can be found in Baddeley et al., (2017), Anderes et al., (2020), Cronie et al., (2020), Moradi and Mateu, (2020), Baddeley et al., (2021), Porcu et al., (2023), Bolin et al., (2024), Møller and Rasmussen, (2024), Moradi and Sharifi, (2024), Filosi et al., (2025), González-Pérez et al., (2025), Bolin et al., (2025), and Alegría et al., (2026).
@@
A second-order Gaussian random field (GRF), namely $\{X(s):s\in\mathcal{L}\}$, is completely characterized by its mean function, which we assume to be zero for simplicity, and by its covariance function, given by
$K(s,t)={\rm cov}[X(s),X(t)],\qquad s,t\in\mathcal{L}.$
The construction of valid and versatile covariance functions is challenging because of the requirement that the covariance matrix of $(X(s_{1}),\ldots,X(s_{N}))^{\top}$, given by
@@
must be positive semidefinite, for all $N\in\mathbb{N}$ and for all collection of spatial locations $s_{1},\dots,s_{N}$ in $\mathcal{L}$. In other words, $K$ must be a positive semidefinite function. The spectral theory of random fields on Euclidean spaces is useful to check for positive semidefiniteness of a candidate function (see, e.g., Yaglom,, 1987). In the linear network setting, however, analogous characterizations are unavailable.
@@
with $\text{d}(\cdot,\cdot)$ standing for a distance metric on $\mathcal{L}$. In practice, treating network coordinates as ordinary Euclidean coordinates can distort the proximity between locations and produce unrealistic results. To address this, Anderes et al., (2020) established positive definiteness for certain classes of isotropic covariance functions under the resistance metric (an extension of the resistance metric from electrical network theory) and showed that, under the geodesic (shortest-path) metric, these classes are valid only under restrictive assumptions on the network structure. Both the resistance and geodesic metrics, unlike the Euclidean metric, take the network topology into account.
@@
Although isotropy is a convenient assumption, it can be overly restrictive. In geostatistical applications, there is increasing demand for more flexible models, particularly when analyzing data over large domains where the properties of the process may vary across space. See Bolin et al., (2025) for a recent development of non-stationary random fields on metric graphs formulated as solutions to stochastic partial differential equations, and an application to traffic speed data. In the context of log Gaussian Cox processes (LGCPs) (Cox,, 1955), where the intensity of point occurrences is governed by a log Gaussian random field, isotropy of the underlying GRF prevents the number of points and the spatial extent of point clusters from varying across the domain. This limitation has been studied for point processes in Euclidean spaces by Dvořák et al., (2019). Similar issues arise on linear networks, as discussed by Baddeley et al., (2017), encouraging the development of more flexible models.
@@
Motivated by this, we propose a general class of non-stationary covariance functions on linear networks, that is, covariance functions not solely determined by distance. The proposed covariance models allow both the variance and the correlation range of the random field to vary across the network, providing a network analogue of models previously proposed in Euclidean spaces (Paciorek,, 2003; Paciorek and Schervish,, 2006; Anderes and Stein,, 2011; Kleiber and Nychka,, 2012; Mateu et al.,, 2013; Liang and Marcotte,, 2016). Our framework produces covariance functions that are locally isotropic with respect to the resistance metric, ensuring interpretability and alignment with the topology of the network.
@@
We exploit the connection between the proposed covariance functions and the class of generalized Stieltjes functions (see Koumandos and Pedersen,, 2019; Menegatto,, 2020 for a detailed treatment) to provide examples of non-stationary covariance functions in closed form.
The proposed covariance functions admit explicit stochastic representations of the associated random fields. We discuss the importance of these representations in the design of computationally efficient geostatistical simulation algorithms. In addition, the framework is extended to construct matrix-valued covariance functions for vector-valued GRFs on linear networks, a relevant setting for analyzing multiple georeferenced variables simultaneously.
@@
To address statistical inference, we revisit the weighted local likelihood approach, originally introduced by Anderes and Stein, (2011) for Euclidean spaces, and adapt it to the network setting. We then conduct a simulation study using synthetic data generated on the street network of the University of Chicago neighborhood, in which we estimate the local correlation range, examine the behavior of the method under different weighting schemes, and evaluate its computational performance.
@@
The paper is organized as follows. Section 2 reviews the definition and properties of the resistance metric. Section 3 presents the main theorem, examples, and illustrations. In Section 4 we provide a stochastic representation and extend the framework to construct matrix-valued covariance functions. Section 5 describes the weighted local likelihood method and reports the results of the simulation study. Section 6 concludes the paper with a discussion and directions for future research. Proofs are provided in the appendices.
@@ 2 Resistance metric
This section reviews the definition and properties of the resistance metric introduced in Anderes et al., (2020), as it is essential for building the non-stationary covariance functions proposed in this work.
We begin by representing the linear network as a metric graph $(\mathcal{V},\mathcal{E})$, where the vertex set $\mathcal{V}=\{v_{1},\dots,v_{n}\}$ corresponds to the endpoints of $\mathcal{L}$, and the edge set $\mathcal{E}$ corresponds to the interiors of the linear segments $\ell_{1},\dots,\ell_{m}$. The length of $\ell_{i}$ is denoted by $|\ell_{i}|$. We write $v_{i}\sim v_{j}$ if the distinct vertices $v_{i}$ and $v_{j}$ are connected by an edge. For any pair of vertices $v_{i},v_{j}\in\mathcal{V}$, the conductance function is defined as
@@
where $v^{*}\in\mathcal{V}$ is an arbitrary reference vertex and $c(v_{i})=\sum_{v\sim v_{i}}\text{con}(v_{i},v)$. The resulting matrix can be viewed as a modification of the graph Laplacian, where the inclusion of an additional unit term on the diagonal guarantees strict positive definiteness. This construction is invariant to the choice of $v^{*}$ (Anderes et al.,, 2020).
@@
    where $B_{1},\dots,B_{m}$ are independent standard Brownian bridges, each defined on an interval of length $|\ell_{i}|$, starting and ending at zero, and independent of $A_{0}$. Since each $B_{i}$ is defined on an interval, writing $B_{i}(s)$ implicitly identifies points $s$ along $\ell_{i}$ with points in this interval via arc length parameterization.
For additional details on the construction of the auxiliary GRF, we refer the reader to Anderes et al., (2020) and Møller and Rasmussen, (2024).
@@
The resistance metric on $\mathcal{L}$ can now be defined as the variogram of the auxiliary GRF in Equation (2.1), that is,
$\text{d}_{R}(s,t)={\rm var}[A(s)-A(t)],\qquad s,t\in\mathcal{L}.$ (2.2)
@@
This metric extends the classical resistance metric from electrical network theory (Klein and Randić,, 1993), originally defined on the vertex set, to the continuum of points along the edges of the graph. When evaluated on any additional points along the edges, $\text{d}_{R}$ coincides with the metric obtained from the corresponding discrete electrical network. Interestingly, the resistance and geodesic distances coincide when $\mathcal{L}$ is a tree, that is, a connected graph with no cycles. Møller and Rasmussen, (2024) provide more fully developed algebraic expressions for the resistance metric, which facilitate its implementation.
@@
We turn to a description of the range of isotropic covariance functions attainable using the resistance metric. A function $\varphi:[0,\infty)\to\mathbb{R}$ is said to be completely monotone on $0,\infty)$ if it is continuous on $0,\infty)$, infinitely differentiable on $(0,\infty)$, and satisfies $(-1)^{j}\varphi^{(j)}(z)\geq 0$, $z>0$, for every integer $j\geq 0$, where $\varphi^{(j)}$ denotes the $j$-th derivative of $\varphi$ and $\varphi^{(0)}=\varphi$. Anderes et al., (2020) showed that for any completely monotone function $\varphi$, the function $K(s,t)=\varphi\big(\text{d}_{R}(s,t)\big)$ is (strictly) positive definite on $\mathcal{L}\times\mathcal{L}$. As a consequence, many popular parametric classes of covariance functions developed for Euclidean spaces are also valid with respect to the resistance metric for general linear networks (see Anderes et al.,, 2020, Table 1).
@@
In contrast, the geodesic distance provides a more restrictive alternative. For example, if $\mathcal{L}$ contains three distinct paths connecting two points $s,t\in\mathcal{L}$, the classical exponential covariance function, defined with respect to the geodesic metric, is no longer positive definite on this network (Anderes et al.,, 2020). Consequently, covariance functions defined in terms of completely monotone functions, which are mixtures of exponential covariance functions, are no longer positive definite on such networks when the geodesic metric is used. This considerably restricts the use of the geodesic distance, as many real situations, such as phenomena on street networks, typically do not satisfy this condition.
@@
Møller and Rasmussen, (2024) also emphasize, in the context of point processes on linear networks, the advantages of the resistance metric over the geodesic metric in terms of flexibility. In particular, the use of $\text{d}_{R}$ allows one to model stronger clustering effects in Cox processes and higher degree of repulsiveness in determinantal point processes.
@@ 3.1 Main result
Let $a:\mathcal{L}\rightarrow(0,\infty)$ and $b:\mathcal{L}\rightarrow(0,\infty)$, and define $\alpha(s,t)=[a(s)+a(t)]/2$. Let $\text{d}_{R}(s,t)$ be the resistance metric between $s$ and $t$, and let $F$ denote a positive finite measure on $(0,\infty)$. Then, the function
$K(s,t)=\displaystyle\int_{0}^{\infty}\frac{\sqrt{b(s)b(t)}}{\sqrt{\alpha(s,t)+\text{d}_{R}(s,t)w}}\text{d}F(w),\qquad s,t\in\mathcal{L},$ (3.1)
@@
Our proof of Theorem 3.1, presented in Appendix A, adapts ideas introduced in Paciorek, (2003) for establishing positive definiteness of non-stationary covariance functions in Euclidean spaces. To the best of our knowledge, the present work is the first to extend this approach to linear networks. The fact that the resistance metric is defined as the variogram of a specific auxiliary GRF gives the proof a distinctive character.
@@
In this framework, the function $b$ provides a way to modulate the locally varying variance of the random field, while the function $a$ regulates how quickly the covariance function decreases with distance. In particular, $a$ can be adjusted to influence the locally adaptive correlation range (the distance beyond which correlations fall below $0.05$).